# Exploring the flow harmonic correlations via multi-particle symmetric and asymmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV

### v2

@@ Abstract
We study multi-particle azimuthal correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. We use initial conditions obtained from a Monte-Carlo Glauber model and evolve them within a viscous relativistic hydrodynamics framework that eventually gives way to a transport model in the late hadronic stage of the evolution. We compute the multi-particle symmetric and asymmetric cumulants and present the results for their sensitivity to the shear and bulk viscosities during the hydrodynamic evolution. We also check their sensitivity to resonance decay and hadronic interactions. We show that these observables are more sensitive to the transport properties than the traditional flow observables.
@@ I.2 Asymmetric Cumulants of Flow Amplitudes
In this paper, we have presented the measurement of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV using a hybrid hydrodynamic model. ${\rm SC}(m,n)$ have been computed in previous studies from hydrodynamic and transport models [43, 42, 37, 44, 45, 46, 47], while ${\rm AC}_{2,1}(m,n)$ have been computed in transport models [45, 46]. ALICE collaboration has measured ${\rm AC}_{2,1}(m,n)$ for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [48]. In addition to symmetric cumulants, various other correlators have also been studied [49, 50, 51, 52]. We show the sensitivity of these measurements to the transport properties, such as the shear viscosity to entropy density ratio, $\eta/s$, and the bulk viscosity to entropy density ratio, $\zeta/s$, of the medium.
@@ II Framework
In the following results, some observables will be presented by scaling with the average number of participants, $\langle N_{part}\rangle$. Table 4 lists the values of $\langle N_{part}\rangle$ in different centrality bins, which are determined using the Monte-Carlo Glauber model. The ${\rm SC}(m,n)$ and ${\rm AC}(m,n)$ observables have been computed using particles with transverse momentum ($p_{T}$) ranges of [0.2, 2.0] GeV and [0.01, 3.0] GeV, respectively.
@@
III Results
@@
III.1 Sensitivity to transport coefficients
III.1.1 Four-particle symmetric cumulants
In this section, we present results for ${\rm SC}(m,n)$ from our simulations. To check their sensitivity to the hydro transport coefficients, we compute these observables for the three sets of ensembles mentioned in Table  2. We observe that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by shear and bulk viscosities. In Fig.[1], we show the effect of shear viscosity on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. In $v_{2}\{4\}$, we see a maximum change of around 10% in the mid-central collisions. While in ${\rm SC}(2,3)$, we see a change of around 40-50% throughout the centrality range. We observe a similar dependence on the bulk viscosity of the medium. Thus, these observables are more sensitive to the transport coefficients. Similar to ${\rm SC}(2,3)$, ${\rm SC}(2,4)$ is also affected by the transport coefficients in the same manner. The comparison of the symmetric cumulants with $v_{3}\{4\}$ also reveals a similar difference in their sensitivity to shear and bulk viscosity.
@@ III.1.1 Four-particle symmetric cumulants
Figure 1: Ratios of $v_{2}\{4\}$ and ${\rm SC}(2,3)$, with different hydro parameter sets, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of $\eta/s$ and $\zeta/s$ on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm SC}(2,3)$ using lines and markers. The red line and the cyan band show the effects of $\eta/s$, and the blue line and the magenta band show the effects of $\zeta/s$.
III.1.2 Six-particle asymmetric cumulants
@@
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Fig.[2], we show the effect of shear viscosity on elliptic flow from the four-particle correlations and ${\rm AC}_{2,1}(2,4)$. For ${\rm AC}_{2,1}(2,4)$, we see a change of around 50% throughout the centrality range (note that 0-10% and 10-20% centrality classes have large statistical uncertainties). We observe that the medium’s bulk viscosity also has a similar effect on these observables. Compared to the traditional flow observables, we see that these are more sensitive to the transport coefficients.
Figure 2: Ratios of $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$, with different hydro parameter sets, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of $\eta/s$ and $\zeta/s$ on $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm AC}_{2,1}(2,4)$ using lines and markers. The red line and the cyan band show the effects of $\eta/s$, and the blue line and the magenta band show the effects of $\zeta/s$.
III.2 Sensitivity to hadronic interactions
III.2.1 Four-particle symmetric cumulants
The hadronic afterburner plays an essential role in describing the data collected at RHIC [43]. To check their sensitivity to hadronic interactions and resonance decay, we compute these observables at the end of the three stages of the evolution mentioned in Table  3. We observe that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by resonance decay and enhanced by hadronic interactions. However, the magnitudes of these effects are different for the two observables. In Fig.[3], we show the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. In $v_{2}\{4\}$, we see a change of around 15% due to resonance decay and 5% due to hadronic interactions plus resonance decay in the mid-central collisions. While in ${\rm SC}(2,3)$, we see a change of around 40% due to resonance decay throughout the centrality range and 40-60% due to hadronic interactions plus resonance decay. ${\rm SC}(2,4)$ is also affected in the same manner. Thus, these observables are more sensitive to resonance decay and hadronic interactions. The comparison of the symmetric cumulants with $v_{3}\{4\}$ also reveals similar differences in their sensitivity to these effects.
@@ III.2.1 Four-particle symmetric cumulants
Figure 3: Ratios of $v_{2}\{4\}$ and ${\rm SC}(2,3)$, computed at the three different stages of the evolution, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm SC}(2,3)$ using lines and markers. The red line and the cyan band show the effects of resonance decay, and the blue line and the magenta band show the effects of hadronic interactions.
III.2.2 Six-particle asymmetric cumulants
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Fig.[4], we show the effects of resonance decay and hadronic interactions on elliptic flow from the four-particle correlations and ${\rm AC}_{2,1}(2,4)$. For ${\rm AC}_{2,1}(2,4)$, we see a change of around 40% due to these in the mid-central collisions. Although the mean values are somewhat similar throughout centrality, due to large uncertainties (about 2X), AC results for 0-10% and 10-20% centralities have been excluded for clarity. ${\rm AC}_{2,1}(2,3)$ is also affected in the same manner. Compared to the traditional flow observables, we see that these are more sensitive to resonance decay and hadronic interactions.
@@ III.2.2 Six-particle asymmetric cumulants
Figure 4: Ratios of $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$, computed at the three different stages of the evolution, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm AC}_{2,1}(2,4)$ using lines and markers. The red line and the cyan band show the effects of resonance decay, and the blue line and the magenta band show the effects of hadronic interactions. AC results for 0-10% and 10-20% centralities, with large uncertainties, have been excluded for clarity.
III.3 Centrality dependence of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$
III.3.1 Four-particle symmetric cumulants
@@
Figure 5: Symmetric cumulant vs centrality from four-particle correlations compared with STAR data [70] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm SC}(2,3)$ is consistently negative while as ${\rm SC}(2,4)$ is consistently positive.
The normalized symmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV, computed using Eq.[5] in momentum space and the corresponding equation in coordinate space for the initial state eccentricity (Eq.[11]) evaluated using the Monte Carlo Glauber model, are presented in Fig.[6].
@@
Figure 6: Normalized symmetric cumulant vs centrality from four-particle correlations compared with STAR data [70] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm NSC}(2,3)$ is consistently negative while as ${\rm NSC}(2,4)$ is consistently positive.
III.3.2 Six-particle asymmetric cumulants

### v3

@@ Abstract
We study multi-particle azimuthal correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. We use initial conditions obtained from a Monte-Carlo Glauber model and evolve them within a viscous relativistic hydrodynamics framework that eventually gives way to a transport model in the late hadronic stage of the evolution. We compute the multi-particle symmetric and asymmetric cumulants and present the results for their sensitivity to the shear and bulk viscosities during the hydrodynamic evolution. We also check their sensitivity to resonance decay and hadronic interactions. We demonstrate that while some of these observables are more sensitive to transport properties than traditional flow observables, others are less sensitive, making them suitable for studying different stages of the evolution.
@@ I.2 Asymmetric Cumulants of Flow Amplitudes
In this paper, we have presented the measurement of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV using a hybrid hydrodynamic model. ${\rm SC}(m,n)$ have been computed in previous studies from hydrodynamic and transport models [43, 44, 42, 37, 45, 46, 47, 48], while ${\rm AC}_{2,1}(m,n)$ have been computed in transport models [46, 47]. ALICE collaboration has performed a systemic study of ${\rm SC}(m,n)$ and ${\rm NSC}(m,n)$ [49]. ALICE collaboration has also measured ${\rm AC}_{2,1}(m,n)$ for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [50]. In addition to symmetric cumulants, various other correlators have also been studied [51, 52, 53, 54]. We show the sensitivity of these measurements to the transport properties, such as the shear viscosity to entropy density ratio, $\eta/s$, and the bulk viscosity to entropy density ratio, $\zeta/s$, of the medium.
@@ II Framework
In the following results, some observables will be presented by scaling with the average number of participants, $\langle N_{part}\rangle$. Table 4 lists the values of $\langle N_{part}\rangle$ in different centrality bins, which are determined using the Monte-Carlo Glauber model. The ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ observables have been computed using particles with transverse momentum ($p_{T}$) ranges of [0.2, 2.0] GeV and [0.01, 3.0] GeV, respectively.
@@
III Error Estimation
The bootstrap method is used to efficiently estimate the error on ensemble-level observables using the Monte-Carlo approach, bypassing the complexities of standard error propagation. It involves randomly selecting events with replacement from an original sample to create bootstrap samples, allowing for the calculation of sampling variance.
Let $\hat{O}$ be the estimator of a statistic on which we intend to find the standard error. To estimate the standard error using the bootstrap method, follow these steps:
  • Given a parent sample of size $n$, construct $B$ independent bootstrap samples $X^{*}_{1},X^{*}_{2},\ldots,X^{*}_{B}$, each with $n$ data points randomly drawn with replacement.
  • Evaluate the estimator for each bootstrap sample.
    $\hat{O}^{*}_{b}=\hat{O}(X^{*}_{b}),\hskip 28.45274ptb=1,2,...,B.$
  • The sampling variance of the estimator can then be calculated as follows:
    $Var(\hat{O})=\frac{1}{B-1}\sum_{b=1}^{B}(\hat{O}^{*}_{b}-\bar{\hat{O}}),$
    where, $\bar{\hat{O}}=\frac{1}{B}\sum_{b=1}^{B}\hat{O^{*}_{b}}$.
The optimal value of $B$ for accurate error estimation varies, but generally, larger values yield better estimates.
IV Results
@@
IV.1 Sensitivity to transport coefficients
@@
Figure 1: Ratios of $v_{2}\{4\}$, ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$, with different hydro parameter sets, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of $\eta/s$ and $\zeta/s$ on $v_{2}\{4\}$, ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$. The plotted results show ratios of $v_{2}\{4\}$ using cyan ($\eta/s$) and magenta ($\zeta/s$) bands and ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$ using red (n=3, $\eta/s$), green (n=3, $\zeta/s$), blue (n=4, $\eta/s$) and orange (n=4, $\zeta/s$) bands. In the legends, the left column show the ratios for checking the effect of $\eta/s$ and the right column show the ratios for checking the effect of $\zeta/s$.
IV.1.1 Four-particle symmetric cumulants
In this section, we present the results for ${\rm SC}(m,n)$ and ${\rm NSC}(m,n)$ obtained from our simulations. To evaluate their sensitivity to the hydro transport coefficients, we computed these observables using the three sets of ensembles listed in Table  2.
We found that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by the effects of shear and bulk viscosities. In Fig.[1], we illustrate how shear viscosity influences $v_{2}\{4\}$ and ${\rm SC}(2,3)$. For $v_{2}\{4\}$, the maximum change observed is about 10% in mid-central collisions. In contrast, ${\rm SC}(2,3)$ shows a change of approximately 40-50% across the centrality range. We also noted a similar dependence on the bulk viscosity of the medium, indicating that these observables are particularly sensitive to the transport coefficients.
The transport coefficients have more significant effects on ${\rm SC}(2,4)$ than on ${\rm SC}(2,3)$. A comparison of symmetric cumulants with $v_{3}\{4\}$ reveals a similar difference in their sensitivity to shear and bulk viscosities. It should be noted that $v_{n}\{4\}$ is more sensitive to the transport coefficients as $n$ increases [72]. This large sensitivity of symmetric cumulants primarily arises from the characteristics of anisotropic flow, as the symmetric cumulants involve higher powers of flow coefficients.
As discussed in the introduction, normalization removes the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the symmetric cumulant. Therefore, it should also eliminate the incidental sensitivity of ${\rm SC}(m,n)$ to shear and bulk viscosity. So, we also examined the sensitivities of the normalized cumulants. From Fig.[1], we observe that ${\rm NSC}(2,3)$ shows little sensitivity to shear viscosity. However, ${\rm NSC}(2,4)$ is more sensitive to both shear and bulk viscosities than $v_{2}\{4\}$.
IV.1.2 Six-particle asymmetric cumulants
@@
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Fig. 1, we illustrate the effect of shear viscosity on $v_{2}\{4\}$ and ${\rm AC}_{2,1}(m,n)$. For ${\rm AC}_{2,1}(2,4)$, we observe a change of approximately 50% across the centrality range, although it is important to note that the 0-10% and 10-20% centrality classes contain large statistical uncertainties. We also find that the medium’s bulk viscosity similarly influences these observables.
The effects of the transport coefficients is found to be greater on ${\rm AC}_{2,1}(2,4)$ than on ${\rm AC}_{2,1}(2,3)$. Compared to traditional flow observables, these observables are more sensitive to transport coefficients. This magnified of asymmetric cumulants sensitivity primarily stems from the characteristics of anisotropic flow, as the asymmetric cumulants involve higher powers of flow.
As discussed in the introduction, normalization removes the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the asymmetric cumulant. Therefore, it should also eliminate the incidental sensitivity of ${\rm AC}_{2,1}(m,n)$ to shear and bulk viscosity. So, we also examined the sensitivities of the normalized asymmetric cumulants. From Fig.[1], it is evident that ${\rm NAC}_{2,1}(2,3)$ shows little sensitivity to shear viscosity. In contrast, ${\rm NAC}_{2,1}(2,4)$ demonstrates greater sensitivity to both shear and bulk viscosities compared to $v_{2}\{4\}$.
IV.2 Sensitivity to hadronic interactions
@@
Figure 2: Ratios of $v_{2}\{4\}$, ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$, computed at the three different stages of the evolution, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$, ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$. The plotted results show ratios of $v_{2}\{4\}$ using cyan (resonance decay) and magenta (hadronic interactions) bands and ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$ using red (n=3, resonance decay), green (n=3, hadronic interactions), blue (n=4, resonance decay) and orange (n=4, hadronic interactions) bands. In the legends, the left column show the ratios for checking the effect of resonance decay and the right column show the ratios for checking the effect of hadronic interactions.
@@
Figure 3: Ratios of $v_{2}\{4\}$, ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$, computed at the three different stages of the evolution, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$, ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$. The plotted results show ratios of $v_{2}\{4\}$ using cyan (resonance decay) and magenta (hadronic interactions) bands and ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$ using red (n=3, resonance decay), green (n=3, hadronic interactions), blue (n=4, resonance decay) and orange (n=4, hadronic interactions) bands. In the legends, the left column show the ratios for checking the effect of resonance decay and the right column show the ratios for checking the effect of hadronic interactions.
Figure 4: Ratios of $v_{2}\{4\}$, ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$, computed at the three different stages of the evolution, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$, ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$. The plotted results show ratios of $v_{2}\{4\}$ using cyan (resonance decay) and magenta (hadronic interactions) bands and ${\rm SC}(2,n)$, ${\rm NSC}(2,n)$, ${\rm AC}_{2,1}(2,n)$, and ${\rm NAC}_{2,1}(2,n)$ using red (n=3, resonance decay), green (n=3, hadronic interactions), blue (n=4, resonance decay) and orange (n=4, hadronic interactions) bands. In the legends, the left column show the ratios for checking the effect of resonance decay and the right column show the ratios for checking the effect of hadronic interactions.
IV.2.1 Four-particle symmetric cumulants
The hadronic afterburner plays a crucial role in interpreting data collected at RHIC [44]. To assess the sensitivity to hadronic interactions and resonance decay, we computed various observables at the end of the three stages of evolution mentioned in Table  3. We found that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by resonance decay and enhanced by hadronic interactions; however, the magnitudes of these effects differ between the two observables.
In Figs. [2][3][4], we illustrate how resonance decay and hadronic interactions impact $v_{2}\{4\}$ and ${\rm SC}(2,3)$. For $v_{2}\{4\}$, we observe a change of about 15% due to resonance decay, with an additional 5% change due to both hadronic interactions and resonance decay in mid-central collisions. In contrast, ${\rm SC}(2,3)$ shows a change of approximately 40% due to resonance decay across the entire centrality range and a change of 40-60% due to the combined effects of hadronic interactions and resonance decay. ${\rm SC}(2,4)$ is affected more strongly by resonance decay and late-stage hadronic interactions than ${\rm SC}(2,3)$. These observables, therefore, demonstrate higher sensitivity to resonance decay and hadronic interactions. The comparison of the symmetric cumulants with $v_{3}\{4\}$ also indicates similar differences in sensitivity to these effects. This increased sensitivity of symmetric cumulants arises mainly from anisotropic flow, with symmetric cumulants incorporating higher powers of flow.
As discussed in the introduction, normalization removes the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the symmetric cumulant. Therefore, it should also eliminate the incidental sensitivity of ${\rm SC}(m,n)$ to resonance decay and late-stage hadronic interactions. So, we also examined the sensitivities of the normalized symmetric cumulants. From Figs. [2][3][4], it is clear that ${\rm NSC}(2,3)$ shows little sensitivity to either resonance decay or late-stage hadronic interactions (UrQMD). In contrast, NSC(2, 4) exhibits greater sensitivity to both resonance decay and late-stage hadronic interactions compared to $v_{2}\{4\}$.
IV.2.2 Six-particle asymmetric cumulants
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Figs.[2][3][4], we show the effects of resonance decay and hadronic interactions on elliptic flow from the four-particle correlations and ${\rm AC}_{2,1}(2,3)$. For ${\rm AC}_{2,1}(2,3)$, we see a change of around 40% due to these in the mid-central collisions. The effects of resonance decay and late-stage hadronic interactions are found to be greater on ${\rm AC}_{2,1}(2,4)$ than on ${\rm AC}_{2,1}(2,3)$. Compared to the traditional flow observables, we see that these are more sensitive to resonance decay and hadronic interactions. This magnified sensitivity of asymmetric cumulants can be attributed to the properties of anisotropic flow, particularly as the asymmetric cumulants incorporate higher powers of flow.
As discussed in the introduction, normalization removes the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the symmetric cumulant. Therefore, it should also eliminate the incidental sensitivity of ${\rm SC}(m,n)$ to resonance decay and late-stage hadronic interactions. So, we also investigated the sensitivities of the normalized asymmetric cumulants. From Figures [2][3][4], it is evident that ${\rm NAC}_{2,1}(2,3)$ shows minimal sensitivity to both resonance decay and late-stage hadronic interactions (as represented by the UrQMD model). In contrast, ${\rm NAC}_{2,1}(2,4)$ demonstrates significantly greater sensitivity to these factors compared to $v_{2}\{4\}$. Therefore, these normalized cumulant measurements have the potential to differentiate between various transport models. Additionally, we examined the sensitivities of the normalized cumulants. From Figs. [2][3][4], it is clear that ${\rm NAC}_{2,1}(2,3)$ shows little sensitivity to either resonance decay or late-stage hadronic interactions (UrQMD). In contrast, ${\rm NAC}_{2,1}(2,4)$ exhibits greater sensitivity to both resonance decay and late-stage hadronic interactions compared to $v_{2}\{4\}$.
As a result, these normalized (a)symmetric cumulant measurements have the potential to differentiate between various models of quark-gluon plasma (QGP) evolution in hydrodynamic and transport models. ${\rm NSC}(2,3)$ and ${\rm NAC}_{2,1}(2,3)$, due to their insensitivity to hydro model parameters, resonance decay, and late-stage hadronic interactions, are particularly useful for constraining the initial state of the system’s evolution.
IV.3 Centrality dependence of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$
IV.3.1 Four-particle symmetric cumulants
@@
Figure 5: (Normalized) Symmetric cumulant vs centrality from four-particle correlations compared with STAR data [73] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm SC}(2,3)$ is consistently negative while as ${\rm SC}(2,4)$ is consistently positive.
The normalized symmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV, computed using Eq.[5] in momentum space and the corresponding equation in coordinate space for the initial state eccentricity (Eq.[11]) evaluated using the Monte Carlo Glauber model, are presented in Fig.[5]. We observe that ${\rm NSC}(2,3)$ shows little sensitivity to shear and bulk viscosities. However, ${\rm NSC}(2,4)$ is more sensitive to both shear and bulk viscosities than $v_{2}\{4\}$. This is due to the fact that $v_{4}$ is the superposition of a nonlinear term that is proportional to $(v_{2})^{2}$ and a linear term $v_{4L}$. This can be expressed as:
$v_{4}=v_{4L}+\chi_{4}(v_{2})^{2},$
where $\chi_{4}$ is the non-linear response coefficient. The term $v_{4L}$ is damped more strongly than the nonlinear term, as noted in Reference [74]. Additionally, $v_{4L}$ is less correlated with $v_{2}$. Consequently, as viscosity increases, the correlation between $v_{4}$ and $v_{2}$ also increases.
@@
IV.3.2 Six-particle asymmetric cumulants